OpenStudy (anonymous):

simplify sqrt(5) * sqrt(10)

OpenStudy (anonymous):

just multiply them together

OpenStudy (anonymous):

\[\sqrt{2} * \sqrt{3} = \sqrt{6}\]

OpenStudy (anonymous):

\[\sqrt{5*10}\] \[\sqrt{50}\] \[5\sqrt{2}\]

OpenStudy (anonymous):

well i was trying to not just give the answer